When we compute a sample standard deviation of a data set, do we subtract the sample mean or the sample median from each of the data values?
Answer #1

Answer :

sample mean

EXPLANATION:

Formula for sample standard deviation (s) is given by:

s=\sqrt{\frac{\sum (x-\bar{x})^{2}}{n-1}},

where \bar{x} is the sample mean.